Answer the following questions. (Hint: Apply the accounting equation.)

a. Intel Corporation had assets equal to $64,186 million and liabilities equal to $13,756 million for a recent year-end. What was the total equity for Intel's business at year-end?

$Answer million

b. At the beginning of a recent year, JetBlue Airways Corporation's assets were $7,549 million and its equity was 2,546 million. During the year, assets increased $44 million and liabilities increased $64 million. What was its equity at the end of the year?

$Answer million

c. At the beginning of a recent year, The Walt Disney Company's liabilities equaled $30,864 million. During the year, assets increased by $2,918 million, and year-end assets equaled $72,124 million. Liabilities increased $2,807 million during the year. What were its beginning and ending amounts for equity?

Beginning equity: $Answer million

Ending equity: $Answer million
Answer #1

a) Total equity = Total assets-Total liabilities = 64186-13756 = 50430 Million

b) Total equity = Total assets-Total liabilities = (7549+44)-(5003+64) = 2526 Million

c) Beginning assets = 72124-2918 = 69206

Beginning equity = 69206-30864 = 38342 Million

Ending equity = 72124-33671 = 38453